OpenStudy (anonymous):

When you solve an equation, how do you know when to add and when to subtract? You may include an example if you need to.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Well Is this what you mean? a+10=15 (subtraction) then you'd know because to solve for a you use inverse operations so, 15-10=5 5+10=15 (addition) x-17=23 23+17=40 check: 40-17=23

OpenStudy (anonymous):

yes thanks